OpenStudy (anonymous):

The sum of the digits of a 2-digit number is 13. If the digits are reversed, the new number is 9 more than the original number. Find the original number.

OpenStudy (mertsj):

Let the number be TU or 10T+U. Just as 35 = 3 times 10 + 5 times 1 so TU = 10 times T + 1 times U.

OpenStudy (mertsj):

So now we must reverse the digits and so we get UT which is 10 times U + 1 times T or 10U+T

OpenStudy (mertsj):

We have the original number: 10T+U and the new number: 10U+T

OpenStudy (mertsj):

The problem says the new number (10U+T) is 9 more than the original number (10T+U)

OpenStudy (mertsj):

The new number 10U+T = 9+ 10T+U or 9U-9T = 9 or U-T=1 Also the problem says U+T=13. So add the two equations and get 2U=14 so U = 7 and T = 4 and the original number is 47

OpenStudy (mertsj):

Sorry U= 7 and T = 6 So the original number is 67

myininaya (myininaya):

now 6+7=13 which makes sense

myininaya (myininaya):

good work mert

OpenStudy (mertsj):

ty